# Optimal Transport on Categorical Data for Counterfactuals using Compositional Data

### v1

@@ Abstract
Recently, optimal transport-based approaches have gained attention for deriving counterfactuals, e.g., to quantify algorithmic discrimination. However, in the general multivariate setting, these methods are often opaque and difficult to interpret. To address this, alternative methodologies have been proposed, using causal graphs combined with iterative quantile regressions Plečko and Meinshausen (2020) or sequential transport Fernandes Machado et al. (2025) to examine fairness at the individual level, often referred to as “counterfactual fairness.” Despite these advancements, transporting categorical variables remains a significant challenge in practical applications with real datasets. In this paper, we propose a novel approach to address this issue. Our method involves (1) converting categorical variables into compositional data and (2) transporting these compositions within the probabilistic simplex of $\mathbb{R}^{d}$. We demonstrate the applicability and effectiveness of this approach through an illustration on real-world data, and discuss limitations.
@@ 1.1 Counterfactuals
Counterfactual analysis is an essential method in machine learning, policy evaluation, economics and causal inference. It involves reasoning about “what could have happened” under alternative scenarios, providing insights into causality and decision-making effectiveness. An example could be the concept of counterfactual fairness, as introduced by Kusner et al. (2017), that ensures fairness by evaluating how decisions would change under alternative, counterfactual conditions. Counterfactual fairness focuses on mitigating bias by ensuring that sensitive attributes, such as race, gender, or socioeconomic status, do not unfairly influence outcomes.
@@
In the counterfactual problem, we consider data $\{(s_{i},\mathbf{x}_{i}),i=1,\cdots,n\}$, where $s$ denotes a binary “treatment” (taking values in $\{0,1\}$). With generic notations, the counterfactual version of $(0,\mathbf{x})$ can be constructed as $(1,T^{\star}(\mathbf{x}))$, where $T^{\star}$ is the optimal transport mapping from $\mathbf{X}|S=0$ to $\mathbf{X}|S=1$, as discussed in Black et al. (2020), Charpentier et al. (2023) and De Lara et al. (2024). Unfortunately, this multivariate mapping is usually both complicated to estimate, and hard to interpret. If $\mathbf{x}$ is univariate, it is simply a quantile interpretation: if $x$ is associated to rank probability $u$ within group $s=0$, then its counterfactual version should be associated with the same rank probability in group $s=1$ (mathematically, $T^{\star}=F^{-1}_{1}\circ F_{0}$, where $F_{j}:\mathbb{R}\rightarrow[0,1],j=\{0,1\}$ denotes the cumulative distribution in group $j$, and $F_{j}^{-1}$ is the generalized inverse, i.e., the quantile function). In higher dimensions, one could consider multivariate quantiles, as in Hallin et al. (2021) or Hallin and Konen (2024), but the heuristics is still hard to interpret. Two primary strategies have been proposed in the literature to assess counterfactual fairness using a mutatis mutandis approach. The first, introduced by Plečko and Meinshausen (2020) and Plečko et al. (2024), is grounded in causal graphs (DAGs). In this framework, the outcome $y$ depends on variables $(s,\mathbf{x})$, where the sensitive attribute $s$ “is a source” (a vertex without parents) and $y$ is a “sink” (a vertex without outgoing edges). The second strategy employs optimal transport (OT) to generate counterfactuals Black et al. (2020); Charpentier et al. (2023); De Lara et al. (2024). Recently, Fernandes Machado et al. (2025) unified these approaches by introducing sequential transport aligned with the “topological ordering” of a causal graph.
For instance, when assessing if a predictive model $\widehat{m}(\mathbf{x})$ is gender neutral (considering binary genders for simplicity purposes), we compare the prediction of a woman with features $\mathbf{x}$ and the prediction of a counterfactual version of that woman, had she been a man. The mutatis mutandis counterfactual, unlike the ceteris paribus version, modifies both the sensitive attribute and other characteristics that may be causally influenced by it. For example, if a person’s height is influenced by their sex, the height in the counterfactual world for a female for a different sex should differ from the factual value. In other words, if $x_{1}$ denotes the height, the mutatis mutandis counterfactual version of a female 5’4 tall is not a male 5’4 tall, but probably 5’10. To map the conditional distribution of features $\mathbf{x}$ between the two groups, Optimal Transport techniques are employed, and the mapping function associates each original observation with its counterfactual version. While obtaining counterfactual values for continuous variables using Optimal Transport is feasible, this becomes much more questionable when dealing with categorical variables. What would be the counterfactual version of a female nurse had that person been male, in the context of sexism? Where would a Black person, living in a specific area, reside in the counterfactual scenario had they been White, in the context of racism? Constructing counterfactuals for categorical variables, which may lack a natural order or distance metric, such as occupation or neighbourhood, can be particularly challenging.
@@ 1.2 The Case of Categorical Variables
For absolutely continuous variables, the approaches of Plečko and Meinshausen (2020); Plečko et al. (2024) on the one hand (based on quantile regressions) and Black et al. (2020); Charpentier et al. (2023); De Lara et al. (2024); Fernandes Machado et al. (2025) (based on optimal transport) are quite similar.
@@ 1.3 Agenda
After recalling notations on Optimal Transport in Section 2, we will discuss how to transform categorical variables with $d$ categories into variables taking values in the simplex $\mathcal{S}_{d}$ in $\mathbb{R}^{d}$, i.e., compositional variables, in Section 3. Then, in Section 4, we review the topological and geometrical properties of the simplex in $\mathbb{R}^{d}$ and discuss two methods for transporting distributions within $\mathcal{S}_{d}$. In Section 5, we consider Gaussian optimal transport based on an alternative representation of the probability vector (in the Euclidean space $\mathbb{R}^{d-1}$). In Section 6 we discuss transport and matching directly within the simplex $\mathcal{S}_{d}$ using an appropriate cost function, rather than the standard quadratic cost (Section 6.1 reviews theoretical results on “Dirichlet transport,” while Section 6.2 addresses empirical matching between observations with an appropriate cost function). Finally, Section 7 presents two illustrations using the German Credit and Adult datasets.
@@ 8 Conclusion
In this article, we introduce a novel approach for constructing counterfactuals for categorical data by transforming them into compositional data using a probabilistic classifier. Our approach avoids imposing arbitrary assumptions about label ordering. However, our methodology is not without limitations. Optimal transport computations, particularly on the simplex, can be computationally intensive for large-scale datasets, posing challenges in high-dimensional settings. Additionally, the reliance on a probabilistic classifier in the initial step introduces potential vulnerabilities. Biases may arise if the classifier is poorly calibrated or inaccurate, which would affect the quality of the subsequent analysis.

### v2

@@ Abstract
Recently, optimal transport-based approaches have gained attention for deriving counterfactuals, e.g., to quantify algorithmic discrimination. However, in the general multivariate setting, these methods are often opaque and difficult to interpret. To address this, alternative methodologies have been proposed, using causal graphs combined with iterative quantile regressions (Plečko and Meinshausen, 2020) or sequential transport (Fernandes Machado et al., 2025) to examine fairness at the individual level, often referred to as “counterfactual fairness.” Despite these advancements, transporting categorical variables remains a significant challenge in practical applications with real datasets. In this paper, we propose a novel approach to address this issue. Our method involves (1) converting categorical variables into compositional data and (2) transporting these compositions within the probabilistic simplex of $\mathbb{R}^{d}$. We demonstrate the applicability and effectiveness of this approach through an illustration on real-world data, and discuss limitations.
@@ 1.1 Counterfactuals
Counterfactual analysis, the third level in Pearl (2009)’s causal hierarchy, is widely used in machine learning, policy evaluation, economics and causal inference. It involves reasoning about “what could have happened” under alternative scenarios, providing insights into causality and decision-making effectiveness. An example could be the concept of counterfactual fairness, as introduced by Kusner et al. (2017), that ensures fairness by evaluating how decisions would change under alternative, counterfactual conditions. Counterfactual fairness focuses on mitigating bias by ensuring that sensitive attributes, such as race, gender, or socioeconomic status, do not unfairly influence outcomes.
@@
In the counterfactual problem, we consider data $\{(s_{i},\mathbf{x}_{i}),i=1,\cdots,n\}$, where $s$ denotes a binary “treatment” (taking values in $\{0,1\}$). With generic notations, the counterfactual version of $(0,\mathbf{x})$ can be constructed as $(1,T^{\star}(\mathbf{x}))$, where $T^{\star}$ is the optimal transport (OT) mapping from $\mathbf{X}|S=0$ to $\mathbf{X}|S=1$, as discussed in Black et al. (2020), Charpentier et al. (2023) and De Lara et al. (2024). Unfortunately, this multivariate mapping is usually both complicated to estimate, and hard to interpret. If $\mathbf{x}$ is univariate, it is simply a quantile interpretation: if $x$ is associated to rank probability $u$ within group $s=0$, then its counterfactual version should be associated with the same rank probability in group $s=1$ (mathematically, $T^{\star}=F^{-1}_{1}\circ F_{0}$, where $F_{j}:\mathbb{R}\rightarrow[0,1],j=\{0,1\}$ denotes the cumulative distribution in group $j$, and $F_{j}^{-1}$ is the generalized inverse, i.e., the quantile function). In higher dimensions, one could consider multivariate quantiles, as in Hallin et al. (2021) or Hallin and Konen (2024), but the heuristics is still hard to interpret. While OT-based counterfactual methods have been proposed to assess counterfactual fairness Black et al. (2020); De Lara et al. (2024), an alternative approach introduced by Plečko and Meinshausen (2020) is grounded in causal graphs (DAGs). In this framework, the outcome $y$ depends on variables $(s,\mathbf{x})$, where the sensitive attribute $s$ “is a source” (a vertex without parents) and $y$ is a “sink” (a vertex without outgoing edges). Recently, Fernandes Machado et al. (2025) unified these approaches by introducing sequential transport aligned with the “topological ordering” of a DAG.
For example, to test whether a predictor $\widehat{m}(\mathbf{x})$ is gender-neutral; let the sensitive attribute $s$ be gender (binary genders for simplicity); compare its output on a woman’s features $\mathbf{x}$ with that on her mutatis mutandis male counterpart. Unlike a ceteris paribus change, which flips $s$ while holding all other features fixed, a mutatis mutandis intervention also adjusts any $x_{j}$ causally influenced by $s$. Thus, if $x_{1}$ is height, the counterfactual of a 5’4" woman would not be a 5’4" man but, say, a 5’10" man, via an OT map. While OT handles continuous attributes naturally, categorical features (e.g. occupation or neighbourhood) lack a canonical distance. As a result, generating counterfactuals (e.g. the male counterpart of a female nurse, or where a Black resident of X would live if they were White) becomes particularly challenging.
@@ 1.2 The Case of Categorical Variables
For absolutely continuous variables, the approaches of Plečko and Meinshausen (2020); Plečko et al. (2024) on the one hand (based on quantile regressions) and Black et al. (2020); Charpentier et al. (2023); De Lara et al. (2024); Fernandes Machado et al. (2025) (based on OT) are quite similar.
@@
While motivated by counterfactual fairness, the primary aim of this study is to present the core of a method for deriving counterfactuals for categorical data, applicable to any context requiring counterfactual analysis. Here, for simplicity, we have set aside considerations related to the assumption of a known Structural Causal Model (SCM). [^1] Details on how the method can be integrated within an SCM are discussed in Appendix C.
@@ 1.3 Agenda
After recalling notations on OT in Section 2, we discuss how to transform categorical variables with $d$ categories into variables taking values in the simplex $\mathcal{S}_{d}$ in $\mathbb{R}^{d}$, i.e., compositional variables, in Section 3. In Section 4, we review the topological and geometrical properties of the probability simplex $\mathcal{S}_{d}\subset\mathbb{R}^{d}$. In Section 5, we introduce the first methodology, which transports distributions within $\mathcal{S}_{d}$ via Gaussian OT. This approach relies on an alternative representation of probability vectors in the Euclidean space $\mathbb{R}^{d-1}$ and assumes approximate normality in the transformed space. In Section 6, we present a second methodology, which operates directly on $\mathcal{S}_{d}$ using a tailored cost function instead of the standard quadratic cost. Theoretical aspects of this “Dirichlet transport” framework are discussed in Section 6.1, while empirical strategies for matching categorical observations are developed in Section 6.2. Section 7 provides two empirical illustrations using the German Credit and Adult datasets.
@@ 8 Conclusion
In this article, we introduce a novel approach for constructing counterfactuals for categorical data by transforming them into compositional data using a probabilistic classifier. Our approach avoids imposing arbitrary assumptions about label ordering. However, our methodology is not without limitations. OT computations, particularly on the simplex, can be computationally intensive for large-scale datasets, posing challenges in high-dimensional settings. Additionally, the reliance on a probabilistic classifier in the initial step introduces potential vulnerabilities. Biases may arise from a poorly calibrated or inaccurate classifier, impacting the quality of the subsequent analysis—especially with scarce categories that may need grouping to apply the proposed method.
Appendix A Complexity of the Main Algorithms
The complexity of Algorithm 2 is $\mathcal{O}(d^{3}\!+\!(n_{0}\!+\!n_{1})d^{2})$ primarily arising from the computation of the class-wise covariance matrices $S_{0}$, $S_{1}$, and the transformation matrix $A$. When the number of classes $d$ becomes large, this cost becomes prohibitive; in such scenarios, alternative OT formulations (beyond the Gaussian OT mapping) are advisable.
Alternatively, Algorithm 3 provides a more scalable approach in high-dimensional settings. Here, $d$ only affects the cost matrix computation with complexity $\mathcal{O}(n_{0}n_{1}d)$, while the dominant computational burden lies in solving the OT problem, which scales as $\mathcal{O}((n_{0}n_{1})^{3/2})$ when using the Operator Splitting Quadratic Program (OSQP) solver.
Appendix B Choice of the Cost Function in Algorithm 3
Section 6 introduces an approach to transporting probability measures directly on the simplex $\mathcal{S}_{d}$. More precisely, Section 6.1 defines the existence of an OT map $T^{\star}$ from a source measure $\mathbb{P}_{0}$ on $\mathcal{S}_{d}$ to a target measure $\mathbb{P}_{1}$ on $\mathcal{S}_{d}$, using the cross-entropy as the cost function, which is referred to as “Dirichlet transport” (Baxendale and Wong, 2022). While this cost function guarantees the theoretical existence of an optimal map, no closed-form expression for $T^{\star}$ is available in practice. Instead, numerical optimization is performed via the Kantorovich formulation of the OT problem to match individuals from $\mathbb{P}_{0}$ to $\mathbb{P}_{1}$ on $\mathcal{S}_{d}$, as described in Section 6.2. Although Euclidean costs, such as the Wasserstein-1 or Wasserstein-2 distances, could be considered, they violate the geometry of the simplex by relying on absolute differences, unlike the cross-entropy, which accounts for relative proportions.
Appendix C Integration of the methodology into a SCM for counterfactual fairness assessment
In Section 3, to transform a categorical variable $x$ into a numerical one, we suggest using a probabilistic classifier based on the features $\mathcal{X}_{-x}$, i.e., all features except the categorical variable. In this appendix, we consider a more global approach, within an SCM, which requires positing a structural causal model in advance. Consider the DAG shown in Figure 8, whose topological order is
$S\;\rightarrow\;X_{1}\;\rightarrow\;X_{2}\;\rightarrow\;X_{3}\;\rightarrow\;Y,$
where $S\in\{0,1\}$ is the sensitive treatment (for example, binary gender), $X_{1}\in\mathbb{R}$ is a numeric feature, $X_{2}\in[\![d_{2}]\!]$ and $X_{3}\in[\![d_{3}]\!]$ are categorical features with $d_{2}$ and $d_{3}$ categories, respectively, and $Y$ is the outcome sink. To generate a counterfactual for an individual with $S=0$, we first flip the treatment by setting $S=1$. We then obtain the numeric counterfactual $X_{1}^{\star}$ using a chosen mapping mechanism (e.g., using optimal transport).
Subsequently, we construct each categorical counterfactual in sequence. For $X_{2}$, we fit a classifier on its parents $(S,X_{1})$ to estimate the conditional distribution $\widehat{P}(X_{2}\mid S,X_{1})$. This yields probability vectors in the simplex $\mathcal{S}_{d_{2}}$ for both the factual group $(S=0)$ and the counterfactual group $(S=1)$, which needs to be converted to a single label $X_{2}^{\star}$ in $\{1,\dots,d_{2}\}$. We suggest either sampling according to the probabilities or selecting the highest-probability category (top-label). We repeat the same process for $X_{3}$, estimating $\widehat{P}(X_{3}\mid S,X_{2})$ and mapping its output into $\{1,\dots,d_{3}\}$. Finally, we predict the counterfactual outcome $Y^{\star}$ with a model conditioned on $(S,X_{1}^{\star},X_{2}^{\star},X_{3}^{\star})$. This sequential, topologically ordered procedure embeds each probabilistic result into its required discrete space, ensuring consistency with the a priori SCM.
[…]
Figure 8: Example of a Structural Causal Model with a sensitive attribute $S$, a numeric variable $X_{1}$, two categorical variables $X_{2}$ and $X_{3}$, and an output variable $Y$.
Matching individuals using their probability vectors on $\mathcal{S}_{d}$ allows to uniquely determine them as it involves continuous distributions on $[0,1]$ for each category. In contrast, if counterfactuals were computed directly from categorical data, one would need to rely on the non-deterministic counterfactual framework described in (Pearl et al., 2016, Chap. 4).